# Learning-based Distributed Model Predictive Control using Multi-Agent Bayesian Optimization

### v1

@@ III Fusion of DMPC and MDP
Let $\boldsymbol{\mathrm{s}}_{i}\in\mathcal{S}_{i}$ and $\boldsymbol{\mathrm{a}}_{i}\in\mathcal{A}_{i}$ denote, respectively, the state and action assigned to the agent $i$ where $\mathcal{S}_{i}$ is the local state space and $\mathcal{A}_{i}$ is the local action space. We further denote $\boldsymbol{\mathrm{s}}=\mathrm{col}\left\{\boldsymbol{\mathrm{s}}_{1},\cdots,\boldsymbol{\mathrm{s}}_{m}\right\}\in\mathcal{S}$ as the state of the multi-agent system. Similarly, the action of the whole system is defined as $\boldsymbol{\mathrm{a}}=\mathrm{col}\left\{\boldsymbol{\mathrm{a}}_{1},\cdots,\boldsymbol{\mathrm{a}}_{m}\right\}\in\mathcal{A}$ while the joint (global) policy is $\boldsymbol{\mathrm{\pi}}=\mathrm{col}\left\{\boldsymbol{\mathrm{\pi}}_{1},\cdots,\boldsymbol{\mathrm{\pi}}_{m}\right\}$. The policy of agent $i$ reads as $\boldsymbol{\mathrm{\pi}}_{i}$. Let us consider an agent acting in a discounted Markov Decision Process (MDP), which is defined by a tuple $\left\{\mathcal{S}_{i},\mathcal{A}_{i},L_{i},\mathbb{P}_{i},\gamma\right\}$ where $L_{i}(\boldsymbol{\mathrm{s}}^{k},\boldsymbol{\mathrm{a}}_{i}^{k})$ is the local cost function, $\mathbb{P}_{i}[\boldsymbol{\mathrm{s}}_{i}^{k+1}|\boldsymbol{\mathrm{s}}_{i}^{k},\boldsymbol{\mathrm{a}}_{i}^{k}]$ is the real state transition dynamics and $\gamma\in(0,1]$ is the discount factor.
III-A Multi-agent MDP
In this paper, we consider a class of multi-agent systems in which each agent aims to contribute to the whole system by minimizing a coupling cost $l_{i}^{c}$ while its own task to minimize the local cost $l_{i}$ is preserved. The total cost assigned to each agent then reads as:
$L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{a}}_{i}^{k}\right)=l_{i}\left(\boldsymbol{\mathrm{s}}_{i}^{k},\boldsymbol{\mathrm{a}}_{i}^{k}\right)+l_{i}^{c}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right).$ (8)
where the augmented state $\boldsymbol{\mathrm{\tilde{s}}}^{k}$ is represented as $\boldsymbol{\mathrm{\tilde{s}}}^{k}=\left[\boldsymbol{\mathrm{s}}_{i}^{k},W_{ij}\left(\boldsymbol{\mathrm{s}}_{j}^{k}\right)\right]^{\top}$. Notice that the coupling cost $l_{i}^{c}$ is defined such that the effect of the coupling constraint (1d) on the centralized problem (1) can be captured during the learning process, detailed in Section IV. Let us define the value function of multi-agent system as
@@ III-A Multi-agent MDP
where the global stage cost function $L$ reads as:
@@
The expectation $\mathbb{E}$ is taken over the distribution of the multi-agent Markov chain resulting from the multi-agent system in closed-loop with joint policy ${\boldsymbol{\mathrm{\pi}}}$.
@@
To decompose the multi-agent value function (9), one can consider a linear value-decomposition as follows [21]:
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\sum_{i=1}^{m}V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right),$
@@ III-B Modifying an approximate multi-agent MDP
Considering the value decomposition method in (11), the $N$-step local value function then reads as:
@@
where $\boldsymbol{\mathrm{\tilde{s}}}^{\ell}=\left[\boldsymbol{\mathrm{s}}_{i}^{\ell},W_{ij}\left(\boldsymbol{\mathrm{s}}_{j}^{\ell}\right)\right]^{\top}$. By modification of the local terminal cost $V_{i}^{f}$ and the local stage cost $L_{i}$, we next show that the value function above can capture the local value function $V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ even if it is constructed based on an $N$-step definition and trajectories are imperfectly generated based on the transition models ^$\hat{\mathbb{P}}_{i}[\boldsymbol{\mathrm{\hat{s}}}_{i}^{k+1}|\boldsymbol{\mathrm{s}}_{i}^{k},\boldsymbol{\mathrm{a}}_{i}^{k}],\forall i\in\left\{1,\ldots,m\right\}$.
@@
where $\boldsymbol{\mathrm{\hat{s}}}^{\ell}=\left[\boldsymbol{\mathrm{\hat{s}}}_{i}^{\ell},W_{ij}\left(\boldsymbol{\mathrm{\hat{s}}}_{j}^{\ell}\right)\right]^{\top}$.
@@
Under Assumption 2, one can find the modified local terminal cost $\hat{V}_{i}^{f}$ and modified local stage cost $\hat{L}_{i}$ such that:
@@
delivers the same local value function as one associated to the true multi-agent MDP on $\Xi$ and for any $N$:
@@
Let us choose the modified local terminal and stage costs as follows:
@@
where $\mathcal{V}_{i}^{+}\left(\hat{\boldsymbol{\mathrm{s}}}^{\ell},\boldsymbol{\mathrm{a}}_{i}^{\ell}\right)=\mathbb{E}\left[V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\hat{\boldsymbol{\mathrm{s}}}^{\ell+1}\right)\Bigg{|}\boldsymbol{\mathrm{a}}^{\ell}=\boldsymbol{\mathrm{\pi}}(\boldsymbol{\mathrm{s}}^{\ell})\right]$. Under Assumption 1, the modified local terminal and the stage costs in (16) are retained finite, and using a telescoping sum, the centralized $N$-step value function (14) is rewritten as follows:
$\qquad\mathbb{E}\Bigg{[}\gamma^{N}V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\hat{\boldsymbol{\mathrm{s}}}^{k+N}\right)+V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\hat{\boldsymbol{\mathrm{s}}}^{k}\right)$
@@
The following equality then holds:
@@
$\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\min_{\boldsymbol{\mathrm{\pi}}_{i}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\min_{\boldsymbol{\mathrm{\pi}}_{i}}V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=V_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right),$ (18)
and the local optimal policy then reads as:
@@
$\boldsymbol{\mathrm{\pi}}_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\in\mathrm{arg}\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\min_{\boldsymbol{\mathrm{\pi}}_{i}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right).$ (19)
III-C Parameterized DMPC

### v2

@@ III Fusion of DMPC and MDP
Let $\boldsymbol{\mathrm{s}}_{i}\in\mathcal{S}_{i}$ and $\boldsymbol{\mathrm{a}}_{i}\in\mathcal{A}_{i}$ denote, respectively, the state and action assigned to the agent $i$ where $\mathcal{S}_{i}$ is the local state space and $\mathcal{A}_{i}$ is the local action space. We further denote $\boldsymbol{\mathrm{s}}=\mathrm{col}\left\{\boldsymbol{\mathrm{s}}_{1},\cdots,\boldsymbol{\mathrm{s}}_{m}\right\}\in\mathcal{S}$ as the state of the multi-agent system. Similarly, the action of the whole system is defined as $\boldsymbol{\mathrm{a}}=\mathrm{col}\left\{\boldsymbol{\mathrm{a}}_{1},\cdots,\boldsymbol{\mathrm{a}}_{m}\right\}\in\mathcal{A}$ while the joint (global) policy is $\boldsymbol{\mathrm{\pi}}=\mathrm{col}\left\{\boldsymbol{\mathrm{\pi}}_{1},\cdots,\boldsymbol{\mathrm{\pi}}_{m}\right\}$ with the policy of agent $i$ denoted as $\boldsymbol{\mathrm{\pi}}_{i}$. Let us consider $m$ agents acting in a discounted multi-agent Markov Decision Process (MDP), which is defined by a tuple $\left\{\mathcal{S},\mathcal{A},L,\mathbb{P},\gamma\right\}$ where $L(\boldsymbol{\mathrm{s}}^{k},\boldsymbol{\mathrm{a}}^{k})$ is the global cost function, $\mathbb{P}[\boldsymbol{\mathrm{s}}^{k+1}|\boldsymbol{\mathrm{s}}^{k},\boldsymbol{\mathrm{a}}^{k}]$ is the true transition model, and $\gamma\in(0,1]$ is the discount factor.
III-A Description of the Multi-agent MDP
In this paper, we consider a class of multi-agent systems in which each agent aims to contribute to the whole system by minimizing a coupling cost $l_{i}^{c}$ while its own task to minimize the local cost $l_{i}$ is preserved. The total cost assigned to each agent is then defined as
$L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{a}}_{i}^{k}\right)=l_{i}\left(\boldsymbol{\mathrm{s}}_{i}^{k},\boldsymbol{\mathrm{a}}_{i}^{k}\right)+l_{i}^{c}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right),$ (8)
where the augmented state $\boldsymbol{\mathrm{\tilde{s}}}^{k}$ is defined as $\boldsymbol{\mathrm{\tilde{s}}}^{k}=\left[\boldsymbol{\mathrm{s}}_{i}^{k},W_{ij}\left(\boldsymbol{\mathrm{s}}_{j}^{k}\right)\right]^{\top}$. Notice that the coupling cost $l_{i}^{c}$ is defined such that the effect of the coupling constraint (1d) on the centralized problem (1) can be captured during the learning process, as detailed in Section IV. From the perspective of each agent $i$, an MDP is then defined by a tuple $\left\{\mathcal{S},\mathcal{A}_{i},L_{i},\mathbb{P}_{i},\gamma\right\}$ with the following transition model
$\mathbb{P}_{i}[\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{a}}_{i}]:=\mathbb{P}\Big{[}\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{a}}_{i}^{k},\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\Big{]}.$ (9)
Let us define the value function of the multi-agent system as
@@ III-A Description of the Multi-agent MDP
where the global stage cost function $L$ reads as
@@
The expectation $\mathbb{E}$ in (10) is taken over the distribution of the multi-agent Markov chain resulting from the closed-loop system with the joint policy ${\boldsymbol{\mathrm{\pi}}}$.
@@
To decompose the multi-agent value function (10), one can consider a linear value-decomposition as [22]
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\sum_{i=1}^{m}V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right).$
@@ III-B Modifying an Approximate Multi-agent MDP
Considering the value decomposition in (12), the $N$-step local value function then is
@@
where $\boldsymbol{\mathrm{\tilde{s}}}^{k}=\left[\boldsymbol{\mathrm{s}}_{i}^{k},W_{ij}\left(\boldsymbol{\mathrm{s}}_{j}^{k}\right)\right]^{\top}$. By modifying the local terminal cost $V_{i}^{f}$ and the local stage cost $L_{i}$, we next show that the value function above can capture the local value function $V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ even if it is constructed based on an $N$-step definition, and trajectories are imperfectly generated based on an inaccurate transition model ^$\hat{\mathbb{P}}_{i}[\boldsymbol{\mathrm{\hat{s}}}^{k+1}|\boldsymbol{\mathrm{s}}^{k},\boldsymbol{\mathrm{a}}_{i}^{k}]$.
@@
where $\boldsymbol{\mathrm{\hat{s}}}^{\ell}=\left[\boldsymbol{\mathrm{\hat{s}}}_{i}^{\ell},W_{ij}\left(\boldsymbol{\mathrm{\hat{s}}}_{j}^{\ell}\right)\right]^{\top}$ and $\mathbb{N}_{\geq k}$ denotes the set of natural numbers greater than $k-1$.
@@
Under Assumption 2, one can find a modified local terminal cost $\hat{V}_{i}^{f}$ and modified local stage cost $\hat{L}_{i}$ such that
@@
delivers the same local value function as one associated with the true multi-agent MDP on $\Xi$ and for any $N$:
@@
Let us choose the modified local terminal and stage costs as
@@
where $\mathcal{V}_{i}^{+}\left(\hat{\boldsymbol{\mathrm{s}}}^{\ell},\boldsymbol{\mathrm{a}}_{i}^{\ell}\right)=\mathbb{E}\left[V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\hat{\boldsymbol{\mathrm{s}}}^{\ell+1}\right)\Bigg{|}\boldsymbol{\mathrm{a}}^{\ell}=\boldsymbol{\mathrm{\pi}}(\boldsymbol{\mathrm{s}}^{\ell})\right]$. Under Assumption 1, the modified local terminal and stage costs in (17) are retained finite, and using a telescoping sum, the centralized $N$-step value function (15) is rewritten as
$\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=\mathbb{E}\Bigg{[}\gamma^{N}V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\hat{\boldsymbol{\mathrm{s}}}^{k+N}\right)+V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\hat{\boldsymbol{\mathrm{s}}}^{k}\right)$ (18)
@@
Theorem 2.
For local value functions $\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$ learned on the transition model ^$\hat{\mathbb{P}}_{i}\Big{[}\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i},\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}^{\star}\Big{]}$, the corresponding local policy $\boldsymbol{\mathrm{\pi}}_{i}$ then converges to the optimal local policy $\boldsymbol{\mathrm{\pi}}_{i}^{\star}$.
Proof.
Let us define the Bellman equation as
$\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\right)$ (19)
@@
$\qquad\qquad+\gamma\mathbb{E}_{\hat{\mathbb{P}}_{i}\Big{[}\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i},\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}\Big{]}}\left[\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k+1}\right)\right].$
From the perspective of other agents $\mathcal{M}_{i}$, the optimal version of the above Bellman equation then reads as
$\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\right)$ (20)
@@
$+\gamma\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\mathbb{E}_{\hat{\mathbb{P}}_{i}\Big{[}\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i},\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}\Big{]}}\left[\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k+1}\right)\right].$
By folding $\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}$ into ^$\hat{\mathbb{P}}_{i}$, we then obtain
$\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\right)$ (21)
$\qquad+\gamma\mathbb{E}_{\hat{\mathbb{P}}_{i}\Big{[}\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i},\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}^{\star}\Big{]}}\left[\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k+1}\right)\right].$
We then minimize (21) w.r.t $\boldsymbol{\mathrm{\pi}}_{i}$ as
$\min_{\boldsymbol{\mathrm{\pi}}_{i}}\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\right)$ (22)
$+\gamma\mathbb{E}_{\hat{\mathbb{P}}_{i}\Big{[}\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i},\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}^{\star}\Big{]}}\left[\min_{\boldsymbol{\mathrm{\pi}}_{i}}\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k+1}\right)\right],$
and from Theorem 1, the following equality then holds
$\min_{\boldsymbol{\mathrm{\pi}}_{i}}\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=\min_{\boldsymbol{\mathrm{\pi}}_{i}}\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}V_{i}^{\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=V_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right).$ (23)
Using (23), we then rewrite (22) as
$V_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)=L_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\right)$ (24)
$\qquad\qquad\qquad+\gamma\mathbb{E}_{\hat{\mathbb{P}}_{i}\Big{[}\boldsymbol{\mathrm{\tilde{s}}}^{k+1}|\boldsymbol{\mathrm{\tilde{s}}}^{k},\boldsymbol{\mathrm{\pi}}_{i},\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}^{\star}\Big{]}}\left[V_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k+1}\right)\right].$
Then, if there exists $\bar{\boldsymbol{\mathrm{\pi}}}_{i}\neq\boldsymbol{\mathrm{\pi}}_{i}^{\star}$ such that $\bar{V}_{i}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\leq V_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)$, it will contradict the Bellman equation (24) so that $\boldsymbol{\mathrm{\pi}}^{\star}$ is not the optimal joint policy. We then have that
$\boldsymbol{\mathrm{\pi}}_{i}^{\star}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right)\in\mathrm{arg}\min_{\boldsymbol{\mathrm{\pi}}_{i}}\min_{\boldsymbol{\mathrm{\pi}}_{j\in\mathcal{M}_{i}}}\hat{V}_{i}^{N,\boldsymbol{\mathrm{\pi}}}\left(\boldsymbol{\mathrm{\tilde{s}}}^{k}\right).$ (25)
III-C Parameterized DMPC for Local Policy Approximation